# Quantum walks, quantum cellular automata, quantum field theory, symmetry, interactions

### v1

@@ Abstract
We show that free QED is equivalent to the continuous-space-and-time limit of Fermi and Bose lattice quantum cellular automata theories derived from quantum random walks satisfying simple symmetry and unitarity conditions. In doing so we define the Fermi and Bose theories in a unified manner using the usual fermion internal space but a boson internal space that is six-dimensional. We show that the reduction to a two-dimensional boson internal space (two helicity states arising from spin-1 plus the photon transversality condition) comes from restricting the quantum cellular automaton theory to positive energies. We briefly examine common symmetries of quantum cellular automata, and how time-reversal symmetry demands the existence of negative-energy solutions. These solutions produce a tension in coupling the Fermi and Bose theories, in which the strong locality of quantum cellular automata seems to require a nonzero amplitude to produce negative-energy states, leading to an unphysical cascade of negative-energy particles. However, we show in a 1D model that by extending interactions over a larger (but finite) range it is possible to exponentially suppress the production of negative-energy particles to the point where they can be neglected.
@@
Quantum walks, quantum cellular automata, quantum field theory, symmetry, interactions
@@ 1 Introduction
Quantum walks (QWs) Aharonov93 ; Ambainis01 ; Aharonov01 ; Kempe03 connect information theory with physical law by providing a discrete model of particle dynamics. In such systems, a quantum walk is based on a simple set of principles and symmetries, with a particle located at any of the vertices of a graph or lattice. In each time step, the particle moves to one of the neighboring vertices (connected by an edge). The particle’s time evolution is given by a unitary transformation that, in the continuum limit, can lead to a relativistic wave equation such as the Weyl or Dirac equations, or to Maxwell’s equations BialynickiBirula94 ; Watrous95 ; Meyer96a ; Meyer96 ; Bracken07 ; Chandrashekar10 ; DAriano14 ; Arrighi14 ; Farrelly14 ; MlodinowBrun18 . In recent years this correspondence has been extended to quantum field theories (QFTs) Bisio15a ; Bisio15b ; Mallick16 ; MlodinowBrun20 ; BrunMlodinow20 ; Arrighi20 ; MlodinowBrun21 . The idea is that one can construct a quantum cellular automaton (QCA) from the quantum walk and then recover the desired Lorentz invariant QFT in the limit of continuous time and energies.
@@
Quantum information processing comprises the action of a sequence of unitary operations (quantum gates) on an initial state of a set of quantum systems (e.g., qubits, which are the fundamental units of quantum information). In QFT the time development of a quantum field is given by the action of a unitary operator on a state describing quantum particles or the creation and annihilation operators corresponding to them. The QCA/QFT correspondence described here suggests that systems of the latter type can arise from systems of the former type. This reveals a deep and fundamental connection between those two seemingly unrelated systems.
@@
For example, quantum field theories are traditionally constructed by starting with a classical field and then quantizing it, while the QCA/QFT correspondence suggests that what is fundamental is not a classical field but an underlying quantum walk. This is reminiscent of Wheeler’s suggestion “It from Bit,” in which he proposed that information plays a significant role in the foundation of physics Wheeler89 . The universe, he suggested, may be fundamentally an information processing system from which the apparent reality of matter somehow emerges. The correspondence may also have more practical implications—providing a new calculational approach to simulation on a quantum computer Farrelly20 ; Eon23 ; Lloyd96 .
@@
But there is another reason to consider the relation of discrete models to quantum field theory: the possibility that space may indeed be discrete, or at least that there is different physics at the Planck scale. Though continuous spacetime is an almost universal assumption in physics, even in classical physics the idea of a point particle such as an electron inevitably leads to divergences. Gravity, too, does not allow an arbitrarily high concentration of gravitational charge (i.e., energy) in a small region of spacetime, because if too concentrated an object will collapse to a black hole. Yet to make observations at shorter and shorter distances, one must focus ever larger energies in a small volume. In both QFT and general relativity (GR) these short-range problems become serious—both theories are plagued by singularities and infinities at short distances. In QFT the usual remedy is renormalization but the methods are mathematically dubious. In GR, it is broadly accepted that at small distance scales the theory breaks down and is no longer applicable. As a result, some have speculated that a quantum theory of gravity will naturally lead to the idea that spacetime is not continuous, with the likely values of a fundamental unit of space and time being (at least roughly) the Planck length $\ell\approx 10^{-33}$ cm and the Planck time $t\approx 10^{-43}$ sec. In this view, spacetime is composed of discrete, tiny cells, as is the case in the QCA models of field theory.
@@
An obvious objection to discrete spacetime is that we would lose Lorentz invariance. But though the lattice QCA models do not admit Lorentz invariance, it is recovered in the limiting case of small lattice spacing (or long wavelength). Lorentz symmetry, then, could be an approximation, and effective symmetry valid at length scales long compared to the fundamental length. Probing shorter lengths would uncover small deviations from the Dirac evolution. In particular, a lattice structure would be reflected in a modification of the free-particle dispersion relation. In a discrete universe with a fundamental length and time on the Planck scale, current technology could not distinguish between full Lorentz invariance and the approximate invariance of the discrete system. This situation may soon change and experiments involving neutron diffraction have been proposed to detect the Lorentz non-invariance that would result if a lattice-based theory represented the true laws of nature, with the Dirac theory being only a continuous approximation BrunMlodinow19 .
@@
In this paper we will show how free Quantum Electrodynamics (QED) arises from quantum walks on a lattice and their corresponding QCAs. We will detail how imposing a handful of basic principles leads to the Fermi (Dirac) and Bose (Maxwell) theories and discuss how they can be coupled. We show that a difficulty in going to an interacting theory is that the symmetries of these QCAs—particularly translation and time-reversal symmetry—lead to theories with both positive- and negative-energy solutions. These solutions are highly delocalized, which means that localized couplings will in general lead to the production of negative-energy particles. We discuss this problem, analyze a one-dimensional example, and look at ways to overcome it.
@@
In Section 2 we begin by defining appropriate discrete-time quantum walks for fermions and bosons on a lattice MlodinowBrun18 ; MlodinowBrun21 . The walk represents the one-particle sector of the theory. Once this is defined, it can be promoted to the multi-particle case in much the same way that one defines a Fock space from a single-particle theory. In Section 3 we impose the condition that the walk not have a preferred lattice axis. We’ll see this implies that although the dynamics will not be Lorentz invariant for any finite lattice spacing, Lorentz invariance will be recovered in the long wavelength limit. In Section 4 we diagonalize the physical space (but not the internal space) factor in the evolution operator by changing to the momentum representation of the wave function. For low-momentum fermions with mass, and bosons that are massless, we take the interval $\Delta t$ to be small and define an approximate time derivative operator, allowing us to obtain the momentum space representation of the Dirac and Maxwell equations. In Section 5 we show it is possible to promote the single-particle fermionic and bosonic theories to multi-particle QCAs that reduce to the usual Dirac and Maxwell field theories in the long-wavelength limit. In Section 6 we discuss symmetries of the lattice theories, and their roles, especially the relation of time reversal symmetry to negative energy states. In Section 7 we discuss issues in the inclusion of interactions between the Fermi and Bose theories. Section 8 presents a short summary of our results.
@@ 2 Structure of the walks
In deriving quantum field theories of fermions and bosons from quantum cellular automata, the first step is to define an appropriate discrete-time quantum walk on a lattice BialynickiBirula94 ; Bracken07 ; Chandrashekar10 ; DAriano14 ; Arrighi14 ; Farrelly14 ; MlodinowBrun18 ; MlodinowBrun21 . The walk represents the one-particle sector of the theory. One this is defined, it can be promoted to the multi-particle case in much the same way that one defines a Fock space from a single-particle theory.
The quantum walk is a unitary analog of a classical random walk. In a classical random walk, in each time step the particle has some probability of moving to the neighboring vertices. In a quantum walk, the evolution is unitary, and the particle can move to a superposition of positions.
We can think of the possible particle positions as being the vertices of a graph, with edges between neighboring vertices. For the quantum walks in this paper, the graph is always regular, with every vertex having the same number of neighbors. The number of neighbors is the degree of the graph, labeled by d. We can label each of the $d$ outgoing edges from a vertex $1,\ldots,d$, with that label describing a direction the particle can move. To start with, we will assume that this graph is finite, but we will later consider the infinite limit.
To maintain both unitarity and nontrivial dynamics, the particle generally must have an internal degree of freedom, a “coin” space. The Hilbert space has the form $\mathcal{H}_{\rm QW}=\mathcal{H}_{X}\otimes\mathcal{H}_{C}$, where $\mathcal{H}_{X}$ is the Hilbert space of the particle position, and $\mathcal{H}_{C}$ is the coin space. As we’ll see, for fermions the coin space has even dimension and the coin space operators that appear in the time evolution are operators of half-integral spin. For bosons the corresponding operators are rotation group generators with integral spin.
The lattice we consider is the bcc lattice, consisting of a set of vertices $\mathbf{x}\equiv(x,y,z)=(i\Delta x,j\Delta x,k\Delta x)$, where $i,j,k$ are integers and $\Delta x$ is a fixed lattice spacing. Thus two vertices are neighbors if one is (x,y,z) and the other is $(x\pm\Delta x,y\pm\Delta x,z\pm\Delta x)$. A discrete-time walk on this lattice involves moving one unit $\Delta x$ in each of the three cardinal directions at each time step. In the standard formulation of a quantum walk in three spatial dimensions, the evolution from one time to the next is given by a unitary evolution operator $U_{\rm QW}$ of the form
@@
The $\{S_{j}\}$ are shift operators that move the particle from its current position to its neighbor in the direction $j$ and the operators $A_{j}$ and the unitary $C$ act on an internal space. $C$ scrambles the faces, so that one does not constantly move in the same direction. The overall unitarity of the walk is not guaranteed: unitarity requires that the operators $\{A_{j}\}$, which act on an internal space, satisfy the condition
$UU^{\dagger}=\sum_{j,j^{\prime}}S_{j}S^{\dagger}_{j^{\prime}}\otimes A_{j}A^{\dagger}_{j^{\prime}}=I.$ (2)
In this paper we use a different formulation MlodinowBrun18 ; Chandrashekar11 ; Chandrashekar13 . We decompose each time step into three successive moves, one in each cardinal direction:
@@
In the above, if the particle is a fermion, at each time step it will move to a neighboring position in each direction. If it is a boson, there is the additional possibility that it remains in place when one or two of the direction factors is applied, but not all three (which would correspond to nothing happening, and thus can be excised from our description). Thus, for fermions the $P_{j}^{\pm}$ operators are three pairs of orthogonal projectors acting on the coin space, satisfying $P_{j}^{+}+P_{j}^{-}=I$. For bosons the $P_{j}^{\pm}$ and $P_{j}^{0}$ operators are three triplets of orthogonal projectors acting on the coin space, satisfying $P_{j}^{+}+P_{j}^{-}+P_{j}^{0}=I$.
The unitary coin flip operator is $C(\theta)=e^{-i\theta Q}$ where $Q$ is a Hermitian operator that acts on the internal space. $Q$ flips the forward and backward directions, i.e., $QP_{j}^{+}Q=P_{j}^{-}$. For bosons, we also have $QP_{j}^{0}Q=P_{j}^{0}$. We can, without loss of generality, assume that $Q^{2}=I$ and ${\rm Tr}\{Q\}=0$. Note that the equations we require for $Q$ are obeyed by the parity transformation, $P$: $P{|\mathbf{x}\rangle}={|-\mathbf{x}\rangle}$ implies that $PP_{j}^{+}P=P_{j}^{-}$, $PP_{j}^{-}P=P_{j}^{+}$.
@@
This novel formulation is inherently unitary. For fermions the $\gamma$ matrices (and the Dirac equation as the long wavelength limit) arise from the condition that the different walk axes and directions be uncorrelated (as we explain below). We get an analogous result for bosons, leading to the Maxwell equations.
@@ 3 Internal Space Operators from Symmetry Requirements
The Dirac and Maxwell equations are Lorentz invariant but one cannot demand rotational invariance of a walk on a discrete lattice. Here we impose the condition that the walk not have a preferred lattice axis. We’ll see that although the dynamics will not be Lorentz invariant for any finite lattice spacing, Lorentz invariance will be recovered in the long wavelength limit.
By “not having a preferred axis” we mean that it should be equally likely for the walker to move in any of the eight allowed directions. This is equivalent to demanding that movement along the different axes is, generically, not correlated. A move in the positive $X$ direction, for example, should not bias the move in the positive or negative $Y$ or $Z$ directions. To achieve this, we impose a noncorrelation condition on the 3D walk, which we call the equal norm condition.
@@
If we introduce operators $\Delta P_{j}=P_{j}^{+}-P_{j}^{-}$, we can see that this equal norm condition is equivalent to requiring that the operators, $\Delta P_{j}$ all anti-commute with each other. Given that, it is easy to see that $\{Q,\Delta P_{j}\}=0$ as well. Four mutually anticommuting operators require a coin space of dimension $D\geq 4$, leading us to identify
@@
The equal norm condition has a simple geometric interpretation. There are two projection operators associated with each dimension, corresponding to the forward and backward directions, and they project onto orthogonal spaces. If the internal space is four-dimensional, for example, each of those spaces is two-dimensional, so they project onto orthogonal planes within the four-dimensional internal space. Similarly, the pairs associated with the other two space dimensions project onto orthogonal planes. However, the planes associated with $X$-dimension projection operators are not orthogonal to the planes associated with $Y$- and $Z$-dimension projectors. The equal norm condition requires that all those nonorthogonal planes be oriented at equal angles to each other. It implies that the form of the quantum walk is preserved by rotations that map the lattice axes into each other. Also, if we change the order in which we apply the steps in the $X$, $Y$, and $Z$ directions, we will get another walk of the same form, with the same properties, and this walk will have the same continuum limit.
@@
where again $k=\pm$, $i,j=X,Y,Z$, $i\neq j$, and $c^{\prime}$ is some positive real constant.
Again, there is a simple geometric interpretation, analogous to that in the fermion case. Here, the $P_{j}^{0}$, $P_{j}^{+}$ and $P_{j}^{-}$ are one-dimensional projectors; each projects onto some vector ${|k\rangle}_{j}$, where $k=0,+,-$ and $j=X,Y,Z$. The above conditions tell us:
@@
Note that $J_{i}$ plays the role of $\gamma_{0}\gamma_{i}$ in the above fermion expression for $\Delta P_{X,Y,Z}$. For fermions, in the 1D and 2D cases as well as the massless 3D case, only a two-dimensional internal space is necessary, with $\Delta P_{i}=\sigma_{i}$. For the 3D massive fermion case, the dimension of the internal space must be doubled, with $\Delta P_{i}=\sigma_{i}\rightarrow\Delta P_{i}=\sigma\otimes\sigma_{i}=\gamma_{0}\gamma_{i}$, where the choice of the first $\sigma$ depends on the representation (it is $\sigma_{X}$ for the Dirac representation, with $\gamma_{0}=\sigma_{Z}\otimes 1$). The doubling is due to the number of anticommuting operators required and leads to a theory accommodating two spin states as well as particles of both matter and antimatter.
In our theory the internal space of the boson theory must also be doubled. As we’ll see when we calculate the dynamics, the model just described already includes anti-particle/negative energy states (as well as longitudinal zero-energy states). Since the photon is its own antiparticle the negative energy sector does not carry additional information, and we will restrict ourselves to the positive energy sector of the free theory. We will see however that the dynamics of the above theory ties the polarization to the sign of the energy: right-handed photons come with a positive energy and left-handed photons with a negative energy, and there is no parity operator. To see this last, note that the parity operator $\mathbb{P}$ would satisfy $\mathbb{P}\Delta P_{i}\mathbb{P}=-\Delta P_{i}$, which is not possible for all three operators, i = 1, 2, 3 (see Eq. (12). A doubling of the internal space is thus necessary to allow the inclusion of both transverse polarizations, and a parity operator. We will therefore work with a six-dimensional internal space. Versions of Maxwell’s equations with six-dimensional field vectors have been studied in other contexts (see for example Ref. Mohr ).
To double the internal space we proceed along the lines of the fermion case:
@@
We can now define a parity operator $\mathbb{P}=\sigma_{X}\otimes 1$, which takes $P_{j}^{+}$ into $P_{j}^{-}$ and vice versa. This could be taken as the coin operator $Q=\mathbb{P}$, as we did in the fermion case. However, as we’ll see, the coin generator endows the particle with mass; so, since we are interested in (massless) photons here, we will set $Q=0$.
To preserve a common structure for both fermions and bosons it is useful to define boson gamma matrices in analogy with the usual fermion ones:
$\gamma_{0}^{\rm bose}$$\equiv$ $\sigma_{X}\otimes 1=\mathbb{P},$
@@
4 Momentum Space Picture and the Dirac and Maxwell Equations

### v2

# quantum walks; quantum cellular automata; quantum field theory; symmetry; interactions

@@ Abstract
Recent work has demonstrated a correspondence that bridges quantum information processing and high-energy physics: discrete quantum cellular automata (QCA) can, in the continuum limit, reproduce quantum field theories (QFTs). This QCA/QFT correspondence raises fundamental questions about how matter/energy, information, and the nature of spacetime are related. Here, we show that free QED is equivalent to the continuous-space-and-time limit of Fermi and Bose QCA theories on the cubic lattice derived from quantum random walks satisfying simple symmetry and unitarity conditions. In doing so, we define the Fermi and Bose theories in a unified manner using the usual fermion internal space and a boson internal space that is six-dimensional. We show that the reduction to a two-dimensional boson internal space (two helicity states arising from spin-1 plus the photon transversality condition) comes from restricting the QCA theory to positive energies. We briefly examine common symmetries of QCAs and how time-reversal symmetry demands the existence of negative-energy solutions. These solutions produce a tension in coupling the Fermi and Bose theories, in which the strong locality of QCAs seems to require a non-zero amplitude to produce negative-energy states, leading to an unphysical cascade of negative-energy particles. However, we show in a 1D model that, by extending interactions over a larger (but finite) range, it is possible to exponentially suppress the production of negative-energy particles to the point where they can be neglected.
@@
quantum walks; quantum cellular automata; quantum field theory; symmetry; interactions
@@ 1 Introduction
Quantum walks (QWs) Aharonov93 ; Ambainis01 ; Aharonov01 ; Kempe03 connect information theory with physical law by providing a discrete model of particle dynamics. In such systems, a quantum walk is based on a simple set of principles and symmetries, with a particle located at any of the vertices of a graph or lattice. In each time step, the particle moves to one of the neighboring vertices (connected by an edge). The particle’s time evolution is given by a unitary transformation that, in the continuum limit, can lead to a relativistic wave equation, such as the Weyl or Dirac equation, or to Maxwell’s equations BialynickiBirula94 ; Watrous95 ; Meyer96a ; Meyer96 ; Bracken07 ; Chandrashekar10 ; DAriano14 ; Arrighi14 ; Farrelly14 ; MlodinowBrun18 . In recent years, this correspondence has been extended to quantum field theories (QFTs) Bisio15a ; Bisio15b ; Mallick16 ; MlodinowBrun20 ; BrunMlodinow20 ; Arrighi20 ; MlodinowBrun21 ; Eon23 ; Centofanti24 ; Bisio25 . The idea is that one can construct a quantum cellular automaton (QCA) from the quantum walk and then recover the desired Lorentz-invariant QFT in the limit of continuous time and energies. Several pieces of recent work have grappled with the issues that arise in adding interactions to QCAs Arrighi20 ; MlodinowBrun21 ; Eon23 ; Centofanti24 , which would be necessary to recover interacting QFTs. That is one of the major topics of this paper.
@@
Quantum information processing comprises the action of a sequence of unitary operations (quantum gates) on an initial state of a set of quantum systems (e.g., qubits, which are the fundamental units of quantum information). In QFT, the time development of a quantum field is given by the action of a unitary operator on a state describing quantum particles or the creation and annihilation operators corresponding to them. The QCA/QFT correspondence described here suggests that systems of the latter type can arise from systems of the former type. This reveals a deep and fundamental connection between those two seemingly unrelated systems.
@@
For example, quantum field theories are traditionally constructed by starting with a classical field and then quantizing it, while the QCA/QFT correspondence suggests that what is fundamental is not a classical field but an underlying quantum walk. This is reminiscent of Wheeler’s suggestion “It from Bit,” in which he proposed that information plays a significant role in the foundation of physics Wheeler89 . The universe, he suggested, may be fundamentally an information processing system from which the apparent reality of matter somehow emerges. The correspondence may also have more practical implications—providing a new calculational approach to simulation on a quantum computer Eon23 ; Farrelly20 ; Lloyd96 .
@@
But there is another reason to consider the relation of discrete models to quantum field theory: the possibility that space may indeed be discrete, or at least that there is different physics at the Planck scale. Though continuous spacetime is an almost universal assumption in physics, even in classical physics, the idea of a point particle such as an electron inevitably leads to divergences. Gravity, too, does not allow an arbitrarily high concentration of gravitational charge (i.e., energy) in a small region of spacetime, because if it is too concentrated, an object will collapse to a black hole. Yet, to make observations at shorter and shorter distances, one must focus ever larger energies into a small volume. In both QFT and general relativity (GR), these short-range problems become serious—both theories are plagued by singularities and infinities at short distances. In QFT, the usual remedy is renormalization, but the methods are mathematically dubious. In GR, it is broadly accepted that, at small distance scales, the theory breaks down and is no longer applicable. As a result, some have speculated that a quantum theory of gravity will naturally lead to the idea that spacetime is not continuous, with the likely values of a fundamental unit of space and time being (at least roughly) the Planck length $\ell\approx 10^{-33}$ cm and the Planck time $t\approx 10^{-43}$ s. In this view, spacetime is composed of discrete, tiny cells, as is the case in the QCA models of field theory.
@@
An obvious objection to discrete spacetime is that we would lose Lorentz invariance. But though the lattice QCA models do not admit Lorentz invariance, it is recovered in the limiting case of small lattice spacing (or long wavelength). Lorentz symmetry, then, could be an approximation, an effective symmetry valid at length scales larger than the fundamental length. Probing shorter lengths would uncover small deviations from the Dirac evolution. In particular, a lattice structure would be reflected in a modification of the free-particle dispersion relation. In a discrete universe with a fundamental length and time on the Planck scale, current technology could not distinguish between full Lorentz invariance and the approximate invariance of the discrete system. This situation may soon change, and experiments involving neutron diffraction have been proposed to detect the Lorentz non-invariance that would result if a lattice-based theory represented the true laws of nature, with the Dirac theory being only a continuous approximation BrunMlodinow19 .
@@
In this paper, we will show how free Quantum Electrodynamics (QED) arises from quantum walks on a lattice and their corresponding QCAs. We will detail how imposing a handful of basic principles leads to the Fermi (Dirac) and Bose (Maxwell) theories and discuss how they can be coupled. We show that a difficulty in going to an interacting theory is that the symmetries of these QCAs—particularly translation and time-reversal symmetry—lead to theories with both positive- and negative-energy solutions. These solutions are highly delocalized, which means that localized couplings will in general lead to the production of negative-energy particles. We discuss this problem, analyze a one-dimensional example, and look at ways to overcome it.
@@
In Section 2, we begin by defining appropriate discrete-time quantum walks for fermions and bosons on a lattice MlodinowBrun18 ; MlodinowBrun21 . The walk represents the one-particle sector of the theory. Once this is defined, it can be promoted to the multi-particle case in much the same way that one defines a Fock space from a single-particle theory. In Section 3, we impose the condition that the walk not have a preferred lattice axis. We will see this implies that although the dynamics will not be Lorentz-invariant for any finite lattice spacing, Lorentz invariance will be recovered in the long-wavelength limit. In Section 4, we diagonalize the physical space (but not the internal space) factor in the evolution operator by changing to the momentum representation of the wave function. For low-momentum fermions with mass, and bosons that are massless, we take the interval $\Delta t$ to be small and define an approximate time derivative operator, allowing us to obtain the momentum-space representation of the Dirac and Maxwell equations. In Section 5, we show that it is possible to promote the single-particle fermionic and bosonic theories to multi-particle QCAs that reduce to the usual Dirac and Maxwell field theories in the long-wavelength limit. In Section 6, we discuss symmetries of the lattice theories, along with their roles, especially the relation of time-reversal symmetry to negative-energy states. In Section 7, we discuss issues in the inclusion of interactions between the Fermi and Bose theories. Section 8 presents a short summary of our results.
@@ 2 Structure of the Walks
In deriving quantum field theories of fermions and bosons from quantum cellular automata, the first step is to define an appropriate discrete-time quantum walk on a lattice BialynickiBirula94 ; Bracken07 ; Chandrashekar10 ; DAriano14 ; Arrighi14 ; Farrelly14 ; MlodinowBrun18 ; MlodinowBrun21 . The walk represents the one-particle sector of the theory. Once this is defined, it can be promoted to the multi-particle case in much the same way that one defines a Fock space from a single-particle theory.
The quantum walk is a unitary analog of a classical random walk. In a classical random walk, in each time step, the particle has some probability of moving to the neighboring vertices. In a quantum walk, the evolution is unitary, and the particle can move to a superposition of positions.
We can think of the possible particle positions as being the vertices of a graph, with edges between neighboring vertices. For the quantum walks in this paper, the graph is always regular, with every vertex having the same number of neighbors. The number of neighbors is the degree of the graph, labeled by $d$. We can label each of the $d$ outgoing edges from a vertex $1,\ldots,d$, with that label describing a direction the particle can move in. To start, we will assume that this graph is finite, but we will later consider the infinite limit.
To maintain both unitarity and nontrivial dynamics, the particle generally must have an internal degree of freedom, a “coin” space. The Hilbert space has the form $\mathcal{H}_{\rm QW}=\mathcal{H}_{X}\otimes\mathcal{H}_{C}$, where $\mathcal{H}_{X}$ is the Hilbert space of the particle position, and $\mathcal{H}_{C}$ is the coin space. As we will see, for fermions, the coin space has even dimension, and the coin space operators that appear in the time evolution are operators of half-integer spin. For bosons, the corresponding operators are rotation group generators with integer spin.
The graph we consider is the body-centered cubic lattice, consisting of a set of vertices $\mathbf{x}\equiv(x,y,z)=(i\Delta x,j\Delta x,k\Delta x)$, where $i,j,k$ are integers and $\Delta x$ is a fixed lattice spacing. Thus, two vertices are neighbors if one is (x,y,z) and the other is $(x\pm\Delta x,y\pm\Delta x,z\pm\Delta x)$. A discrete-time walk on this lattice involves moving one unit $\Delta x$ in each of the three cardinal directions at each time step. In the standard formulation of a quantum walk in three spatial dimensions, the evolution from one time to the next is given by a unitary evolution operator $U_{\rm QW}$ BialynickiBirula94 of the form
@@
The $\{S_{j}\}$ are shift operators that move the particle from its current position to its neighbor in the direction $j$, and the operators $A_{j}$ and the unitary $C$ act on an internal space. $C$ scrambles the faces, so that one does not constantly move in the same direction. The overall unitarity of the walk is not guaranteed: unitarity requires that the operators $\{A_{j}\}$, which act on an internal space, satisfy the condition BialynickiBirula94 ; DAriano14
$U_{\rm QW}U_{\rm QW}^{\dagger}=\sum_{j,j^{\prime}}S_{j}S^{\dagger}_{j^{\prime}}\otimes A_{j}A^{\dagger}_{j^{\prime}}=I.$ (2)
In this paper, we use a different formulation MlodinowBrun18 ; Chandrashekar11 ; Chandrashekar13 . We decompose each time step into three successive moves, one in each cardinal direction:
@@
In the above, if the particle is a fermion, at each time step, it will move to a neighboring position in each direction. If it is a boson, there is the additional possibility that it remains in place when one or two of the direction factors are applied, but not all three (which would correspond to nothing happening and thus can be excised from our description). Thus, for fermions, the $P_{j}^{\pm}$ operators are three pairs of orthogonal projectors acting on the coin space, each satisfying $P_{j}^{+}+P_{j}^{-}=I$. For bosons, the $P_{j}^{\pm}$ and $P_{j}^{0}$ operators are three triplets of orthogonal projectors acting on the coin space, each satisfying $P_{j}^{+}+P_{j}^{-}+P_{j}^{0}=I$.
The unitary coin flip operator is $C(\theta)=e^{-i\theta Q}$, where $Q$ is a Hermitian operator that acts on the internal space. $Q$ flips the forward and backward directions, i.e., $QP_{j}^{+}Q=P_{j}^{-}$. For bosons, we also have $QP_{j}^{0}Q=P_{j}^{0}$. We can, without loss of generality, assume that $Q^{2}=I$ and ${\rm Tr}\{Q\}=0$. Note that if $U_{\rm QW}$ is parity-invariant, the equations we require for $Q$ are obeyed by the parity transformation, $\mathcal{P}$: $\mathcal{P}{|\mathbf{x}\rangle}={|-\mathbf{x}\rangle}$ implies that $\mathcal{P}P_{j}^{+}\mathcal{P}=P_{j}^{-}$, $\mathcal{P}P_{j}^{-}\mathcal{P}=P_{j}^{+}$.
@@
This formulation is inherently unitary. For fermions, the $\gamma$ matrices (and the Dirac equation as the long-wavelength limit) arise from the condition that the different walk axes and directions be uncorrelated (as we explain below). We obtain an analogous result for bosons, leading to the Maxwell equations.
@@ 3 Internal Space Operators from Symmetry Requirements
The Dirac and Maxwell equations are Lorentz-invariant, but one cannot demand rotational invariance of a walk on a discrete lattice. Here, we impose the condition that the walk not have a preferred lattice axis. We will see that although the dynamics will not be Lorentz-invariant for any finite lattice spacing, Lorentz invariance will be recovered in the long-wavelength limit.
By “not having a preferred axis”, we mean that it should be equally likely for the walker to move in any of the eight allowed directions. This is equivalent to demanding that movement along the different axes is, generically, not correlated. A move in the positive $X$ direction, for example, should not bias the move in the positive or negative $Y$ or $Z$ directions. To achieve this, we impose a noncorrelation condition on the 3D walk, which we call the equal norm condition.
@@
If we introduce operators the $\Delta P_{j}=P_{j}^{+}-P_{j}^{-}$, we can see that this equal norm condition is equivalent to requiring that the operators $\Delta P_{j}$ all anti-commute with each other. Given that, it is easy to see that $\{Q,\Delta P_{j}\}=0$ as well. Four mutually anti-commuting operators require a coin space of dimension $D\geq 4$, leading us to identify
@@
The equal norm condition has a simple geometric interpretation. There are two projection operators associated with each dimension, corresponding to the forward and backward directions, and they project onto orthogonal spaces. If the internal space is four-dimensional, for example, each of those spaces is two-dimensional, so they project onto orthogonal planes within the four-dimensional internal space. Similarly, the pairs associated with the other two space dimensions project onto orthogonal planes. However, the planes associated with $X$-dimension projection operators are not orthogonal to the planes associated with $Y$- and $Z$-dimension projectors. The equal norm condition requires that all those non-orthogonal planes be oriented at equal angles to each other. It implies that the form of the quantum walk is preserved by rotations that map the lattice axes into each other. Also, if we change the order in which we apply the steps in the $X$, $Y$, and $Z$ directions, we will obtain another walk of the same form, with the same properties, and this walk will have the same continuum limit.
@@
where, again, $k=\pm$, $i,j=X,Y,Z$, $i\neq j$, and $c^{\prime}$ is some positive real constant.
Again, there is a simple geometric interpretation, analogous to that in the fermion case. Here, the $P_{j}^{0}$, $P_{j}^{+}$ and $P_{j}^{-}$ are one-dimensional projectors; each projects onto some vector ${|k\rangle}_{j}$, where $k=0,+,-$ and $j=X,Y,Z$. The above conditions tell us the following:
@@
Note that $J_{i}$ plays the role of $\gamma_{0}\gamma_{i}$ in the above fermion expression for $\Delta P_{X,Y,Z}$. For fermions, in the 1D and 2D cases as well as the massless 3D case, only a two-dimensional internal space is necessary, with $\Delta P_{i}=\sigma_{i}$. For the 3D massive fermion case, the dimension of the internal space must be doubled, with $\Delta P_{i}=\sigma_{i}\rightarrow\Delta P_{i}=\sigma\otimes\sigma_{i}=\gamma_{0}\gamma_{i}$, where the choice of the first $\sigma$ depends on the representation (it is $\sigma_{X}$ for the Dirac representation, with $\gamma_{0}=\sigma_{Z}\otimes 1$). The doubling is due to the number of anti-commuting operators required and leads to a theory accommodating two spin states as well as particles of both matter and antimatter.
In our theory, the internal space of the boson theory must also be doubled. As we will see, when we calculate the dynamics, the model just described already includes anti-particle/negative-energy states (as well as longitudinal zero-energy states). Since the photon is its own anti-particle, the negative-energy sector does not carry additional information, and we will restrict ourselves to the positive-energy sector of the free theory. We will see, however, that the dynamics of the above theory ties the polarization to the sign of the energy: right-handed photons come with a positive energy and left-handed photons with a negative energy, and there is no parity operator. To see the absence of a parity operator, note that the parity operator $\mathcal{P}$ would satisfy $\mathcal{P}\Delta P_{i}\mathcal{P}=-\Delta P_{i}$, which is not possible for all three operators, i = 1, 2, 3 (see Equation (12)). A doubling of the internal space is thus necessary to allow the inclusion of both transverse polarizations, as well as a parity operator. We will therefore work with a six-dimensional internal space. Versions of Maxwell’s equations with six-dimensional field vectors have been studied in other contexts (see, for example, Ref. Mohr ).
To double the internal space, we proceed along the lines of the fermion case:
@@
We can now define a parity operator $\mathcal{P}=\sigma_{X}\otimes 1$, which takes $P_{j}^{+}$ into $P_{j}^{-}$ and vice versa. This could be taken as the coin operator $Q=\mathcal{P}$, as we did in the fermion case. However, as we will see, the coin generator endows the particle with mass; so, since we are interested in (massless) photons here, we will set $Q=0$.
To preserve a common structure for both fermions and bosons, it is useful to define boson gamma matrices in analogy with the usual fermion ones:
$\gamma_{0}^{\rm bose}$$\equiv$ $\sigma_{X}\otimes 1=\mathcal{P},$
@@
We note from the discussion of time-reversal symmetry in Section 6 that the minimum dimension of the internal space that would give two positive-energy solutions is four. However, we know of no construction of a QW with a four-dimensional internal space that gives rise to the Maxwell theory in the long-wavelength limit. So, we proceed with our six-dimensional construction, at the cost of including zero-energy solutions that decouple from the dynamics.
4 Momentum-Space Picture and the Dirac and Maxwell Equations